OpenStudy (anonymous):

I need help with this problem. Can some one help me with this please. Monica Spent 7/9 hours listening too tapes of Beethoven and Brahms. She spent 1/4 hours listening too Beethoven. How many hours were spent listening to brahms.

OpenStudy (anonymous):

is it 1 1/36

OpenStudy (anonymous):

19/36

OpenStudy (anonymous):

7/9-1/4

OpenStudy (anonymous):

Can you type a simplified fractions.

OpenStudy (anonymous):

((7*4)-(9*1) )/4*9

OpenStudy (anonymous):

how would you do that and how?

OpenStudy (anonymous):

together (beethoven and st.. tapes) =7/9 beethoven = 1/4 so st...= 7/9-1/4

OpenStudy (anonymous):

so I subtract 7/9-1/4

OpenStudy (anonymous):

yes

OpenStudy (lstrasmann):

\[\frac{ 7 }{ 9 } - \frac{ 1 }{ 4 }\]You have to find the least common denominator. (LCD) Factors of 9: 9,18,27,36 Factors of 4: 4,8,12,16,20,24,28,32,36 36 is the LCD. Multiply the fractions by 4 and 9 to get denominators of 36. Like this\[\frac{ 7 }{ 9 }\times \frac{ 4 }{ 4 } = \frac{ 28 }{ 36 }\]\[\frac{ 1 }{ 4 }\times \frac{ 9 }{ 9 } = \frac{ 9 }{ 36 }\] \[\frac{ 28 }{ 36 }-\frac{ 9 }{ 36 } = \frac{ 19 }{ 36 }\]

OpenStudy (anonymous):

oh ok. so 19/36 is the answer, because it cant be broken down.. Right

OpenStudy (anonymous):

yes you are correct

OpenStudy (lstrasmann):

Yes.

OpenStudy (anonymous):

does that help

OpenStudy (anonymous):

yes. you guys helped me alot. Thank you so much

OpenStudy (lstrasmann):

You're welcome. :)

OpenStudy (anonymous):

Istrasmann can you help me with another problem please

OpenStudy (lstrasmann):

yes what is it?

OpenStudy (anonymous):

ok. It says add too find the exact answer. Write the answer as a mixed number? 8 7/9 + 6 7/9=

OpenStudy (lstrasmann):

\[14\frac{ 14 }{ 9 }\]which is also\[15\frac{ 5 }{ 9 }\]

OpenStudy (anonymous):

oh ok, 14 14/9 is the answer too that problem.

OpenStudy (anonymous):

How did you get 15 6/9

OpenStudy (lstrasmann):

14/9 is 1 5/9 simplified

OpenStudy (anonymous):

1 5/9 is the answer

OpenStudy (anonymous):

Im getting confused now lol.

OpenStudy (lstrasmann):

no but 14 and 14/9 is improper fraction. so to make it proper, the numerator has to be less than the denominator. so adding 1 5/9 to 14 is 15 5/9

OpenStudy (anonymous):

so its 15 5/9 thats the answer. I think i get it now

OpenStudy (lstrasmann):

yeah 15 5/9 is right

OpenStudy (anonymous):

can you help me with 2 more plems please and Thank you

OpenStudy (lstrasmann):

sure i can.

OpenStudy (anonymous):

A man is 70 4/9inches tall and his son is 63 5/12inches tall. How much taller is the man.. Can you please help me with the step by step process

OpenStudy (lstrasmann):

okay one second i have to check another question i'm overworked lol

OpenStudy (anonymous):

ok

OpenStudy (lstrasmann):

\[70\frac{ 4 }{ 9 } - 63\frac{ 5 }{ 12 }\]the LCD of 9 and 12 is 36. now we need to multiply the fraction parts (not 70 and 63) \[\frac{ 4 }{ 9 }\times \frac{ 4 }{ 4 } = \frac{ 16 }{ 36 }\]\[\frac{ 5 }{ 12 }\times \frac{ 3 }{ 3 }=\frac{ 15 }{ 36 }\] now subtract the fraction then the numbers \[\frac{ 16 }{ 36 }-\frac{ 15 }{ 36 } = \frac{ 1 }{ 36 }\]which cannot be further simplified.\[70-63=7\]so the answer is \[7\frac{ 1 }{ 36 }inches\]

OpenStudy (lstrasmann):

do i get a medal?

OpenStudy (anonymous):

I was jsut about too say 7 1/36 you read my mined

OpenStudy (anonymous):

i figured it out. Im super excited about learning how too do Algebra

OpenStudy (anonymous):

And yes you get a Medal.

OpenStudy (anonymous):

Evalute ( 1/2)^3=

OpenStudy (lstrasmann):

1/2*1/2*1/2=1/8